Consider the system:
{ y=3x+2
{9x-3y=-6

Answer by linh117(31)

a)y=3x+2
b)9x-3y=-6
B)9x-3(3x+2)=-6
9x-9x-6=-6
-18x-6=-6
-18x=0
x=0
A)y=3(0)+2
y=2